Bob left work and traveled to school at an average speed of 25 mph. If it took Bob 35 minutes to reach school, how far is Bob's school from his place of work?
Answer by ewatrrr(24785)

D = rt
D+=+rt%0D%0A25mph*(35/60)hr = 175/12 = 14 7/12 mi